OpenStudy (anonymous):

A triangle has sides of length 5, 8, and 10. What is the measure of its largest angle?

OpenStudy (mertsj):

Use the law of cosines

OpenStudy (anonymous):

I did but can not seem to get the right answer

OpenStudy (mertsj):

\[10^2=5^2+8^2-2(5)(8)\cos A\] \[100-89=80\cos A\] \[\frac{11}{80}=\cos A\]=.1375 \[\cos^{-1} (.1375)=82.1^{o}\]

OpenStudy (anonymous):

That's what I got, but the back of the book said it was 97.9 @Mertsj

myininaya (myininaya):

\[100-89=-80\cos A\]

myininaya (myininaya):

You do get what the book says. Mert just forgot the little negative sign

OpenStudy (anonymous):

How is it the same as what the book says?

myininaya (myininaya):

try using the calculator

myininaya (myininaya):

arccos(-11/80)

OpenStudy (anonymous):

I did so if I use the negative I will get the right answer?

OpenStudy (mertsj):

yes